Brief Exercise 11-07

Crane Company purchased a computer for $8,080 on January 1, 2019. Straight-line depreciation is used, based on a 5-year life and a $1,010 salvage value. On January 1, 2021, the estimates are revised. Crane now feels the computer will be used until December 31, 2022, when it can be sold for $505.

Compute the 2021 depreciation. (Round answer to 0 decimal places, e.g. 45,892.)

Depreciation expense, 2021 $

Entry field with incorrect answer

PLEASE SHOW ALL WORK.
Answer #1

Depreciation for 2021= $2,373.50 or 2374.

Working

Straight line Method for 2019 till 2019 to 2020
A Cost $ 8,080
B Residual Value $ 1,010
C=A - B Depreciable base $ 7,070
D Life [in years left ]                                  5
E=C/D Annual SLM depreciation $               1,414.00

.

Depreciation schedule-Straight line method
Year Book Value Depreciation expense Accumulated Depreciation Ending Book Value
2019 $            8,080.00 $               1,414.00 $                1,414.00 $                 6,666.00
2020 $            6,666.00 $               1,414.00 $                2,828.00 $                 5,252.00

.

Straight line Method - for 2021
A Cost $ 5,252
B Residual Value $ 505
C=A - B Depreciable base $ 4,747
D Life [in years left ] (4-2)                                  2
E=C/D Annual SLM depreciation $               2,373.50